# Asymmetric Bias in Text-to-Image Generation with Adversarial Attacks

### v2

@@ 1 Introduction
    We propose a new metric that is tied to the asymmetric bias of T2I models. This helps us identify vulnerable preconditions and estimate ASR without performing an attack (Figure 0(c)).
@@ 2 Related Works
Adversarial attacks, which perturb inputs to cause models to behave unpredictably, have been a long-studied area in the field of adversarial robustness (Szegedy et al., 2013; Shafahi et al., 2018; Shayegani et al., 2023). Previous studies on adversarial attacks focused on discriminative models involving convolutional neural networks (Athalye et al., 2018; Hendrycks and Dietterich, 2018), while recent work has shifted towards examining generative models such as large language models (LLMs) (Shin et al., 2020; Zou et al., 2023; Liu et al., 2023c; Mo et al., 2023; Cao et al., 2023), vision language models (VLMs) (Dong et al., 2023; Khare et al., 2023; Shayegani et al., 2024), and Text-to-Image (T2I) models detailed below.
@@
Zhuang et al. (2023) were among the first to demonstrate that a mere five-character perturbation could significantly alter the generated images. Tsai et al. (2023) and SneakyPrompt (Yang et al., 2023b) proposed adversarial attacks using genetic algorithms and reinforcement learning algorithms to perturb safe prompts to generate NSFW content. VLAttack (Yin et al., 2023), MMA-Diffusion (Yang et al., 2023a), and INSTRUCTTA (Wang et al., 2023) demonstrated that cross-modality attacks can achieve higher success rates than text-only attacks. For defense, Zhang et al. (2023) proposed Adversarial Prompt Tuning to enhance the adversarial robustness of the image encoder in T2I models.
@@ 3.2 Entity Swapping Dataset
    The input sentence and the target sentence should differ by exactly one noun (i.e., an entity).
@@
    The input sentence and target sentence should be visually distinct.
@@
To ensure that our results were not due to selective data selection, we generated a second dataset of 1,000 pairs deterministically from the test split captions of MS-COCO Lin et al. (2014)  [^3] We released the code to reproduce COCO-Pairs.. We refer to this dataset as COCO-Pairs. Since COCO-Pairs is automatically generated, we attempted to ensure that each data pair satisfies all three requirements. However, generating sentence pairs through stable diffusion and verifying them as visually distinct automatically is not always reliable. We observed some visually non-distinct pairs, such as (“Herd of zebras …”, “Images of zebras …”) within COCO-Pairs despite automatic checks and filtering. See Appendix A for details on dataset curation.
@@ Appendix C Additional Examples of Asymmetric Bias
Sentence Pair (1 / 2) $\boldsymbol{\Delta_{2}}$
ASR $\mathbf{{\color[rgb]{0,0,1}1}\rightarrow{\color[rgb]{1,.5,0}2}}$
ASR $\mathbf{{\color[rgb]{1,.5,0}2}\rightarrow{\color[rgb]{0,0,1}1}}$ Example
@@
a (turtle / fish) swimming in an aquarium. +0.05
0%
90% [Uncaptioned image]
a (robot / human) dancing in the rain. +0.1
0%
60% [Uncaptioned image]
a doodle of a (light bulb / dog) on a blackboard. +0.1
0%
80% [Uncaptioned image]

### v3

@@ 1 Introduction
    We propose a new metric tied to the asymmetric bias of T2I models. This helps us identify vulnerable preconditions and estimate ASR without performing an attack (Figure 1(c)).
@@ 2 Related Works
Adversarial attacks, which perturb inputs to cause models to behave unpredictably, have been a long-studied area in the field of adversarial robustness (Szegedy et al., 2013; Shafahi et al., 2018; Shayegani et al., 2023). Previous studies on adversarial attacks focused on discriminative models involving convolutional neural networks (Athalye et al., 2018; Hendrycks and Dietterich, 2018), while recent work has shifted towards examining generative models such as large language models (LLMs) (Shin et al., 2020; Zou et al., 2023; Liu et al., 2023c; Mo et al., 2023; Cao et al., 2023), Vision Language models (VLMs) (Dong et al., 2023; Khare et al., 2023; Shayegani et al., 2024), and Text-to-Image (T2I) models.
@@
Zhuang et al. (2023) were among the first to demonstrate that a mere five-character perturbation could significantly alter the generated images. Tsai et al. (2023) and SneakyPrompt (Yang et al., 2023b) proposed adversarial attacks using genetic algorithms and reinforcement learning algorithms to perturb safe prompts to generate NSFW content. VLAttack (Yin et al., 2023), MMA-Diffusion (Yang et al., 2023a), and INSTRUCTTA (Wang et al., 2023) demonstrated that cross-modality attacks can achieve higher success rates than text-only attacks. For defense, Zhang et al. (2023) proposed Adversarial Prompt Tuning to enhance the adversarial robustness of VLMs. However, to the best of our knowledge, no comparable defense against non-NSFW attacks exists for T2I models.
@@ 3.2 Entity Swapping Dataset
    The input and target sentences should differ by exactly one noun (i.e., an entity).
@@
    The input and target sentences must be visually distinct.
@@
To ensure that our results were not due to selective data selection, we generated a second dataset of 1,000 pairs deterministically from the test split captions of MS-COCO Lin et al. (2014). We refer to this dataset as COCO-Pairs  [^3] The code to reproduce COCO-Pairs is provided in our codebase.. Since COCO-Pairs is automatically generated, we attempted to ensure that each data pair satisfies all three requirements. However, generating sentence pairs through stable diffusion and verifying them as visually distinct automatically is not always reliable. We observed some visually non-distinct pairs, such as (“Herd of zebras …”, “Images of zebras …”) within COCO-Pairs despite automatic checks and filtering. See Appendix A for details.
@@ Appendix C Additional Examples of Asymmetric Bias
{tblr}
colspec = X[-1,l]X[-1,c]X[-1,c]X[-1,c]X[-1,c,h], Sentence Pair (1 / 2) $\boldsymbol{\Delta_{2}}$ ASR $\mathbf{{\color[rgb]{0,0,1}1}\rightarrow{\color[rgb]{1,.5,0}2}}$ ASR $\mathbf{{\color[rgb]{1,.5,0}2}\rightarrow{\color[rgb]{0,0,1}1}}$ Example
@@
a (turtle / fish) swimming in an aquarium. +0.05 0% 90% [Uncaptioned image]
a (robot / human) dancing in the rain. +0.1 0% 60% [Uncaptioned image]
a doodle of a (light bulb / dog) on a blackboard. +0.1 0% 80% [Uncaptioned image]
@@ G.3 Adversarial Suffixes Do Not Transfer across T2I Models
Table 4 shows that different variants of Stable Diffusion were susceptible to entity-swapping attacks and exhibited similar levels of asymmetric bias on prompt pairs.
Stable Diffusion 1.4 Stable Diffusion 2.1
BSR $\Delta_{2}$ Num. Avg. ASR Num. Avg. ASR
Low Neg. 31 0.119 23 0.174
Low Pos. 26 0.062 19 0.047
High Neg 19 0.553 27 0.6
High Pos. 24 0.25 31 0.171
All All 100 0.218 100 0.264
Table 4: Average ASR of SD 1.4 and SD 2.1 on HQ-Pairs. BSR and $\Delta_{2}$ remain strong predictors in both cases.
However, the adversarial suffixes generated using SD 2.1-base did not work on SD 1.4 and vice versa. SD 2.1-base uses OpenCLIP-ViT/H Cherti et al. (2023) as the text encoder while SD 1.4 uses CLIP ViT-L/14 Radford et al. (2021). Although OpenCLIP-ViT/H and CLIP ViT-L/14 have the same architecture and parameter count, the lack of transferability indicates that training data likely plays the main role in determining adversarial attack success.